# Dumbbell dimer dynamics in three-dimensional chiral fluids

### v1

@@ Introduction
We have shown that the odd viscosity, to leading order, gives rise to the precession dynamics of a dumbbell dimer about the odd viscosity direction $\mathbf{e}$. Here we extend the model to take into account higher order contributions and investigate whether the dynamics also exhibits the out-of-plane dynamics that includes the evolution of $\theta$. By expanding the resistance tensor of a sphere up to the quadratic order in $\lambda$ [45], the velocity equations for the position vectors are
$\dot{\textbf{r}}_{\alpha}=\frac{1}{6\pi\eta a_{\alpha}}\left[\left(1-\frac{\lambda^{2}}{10}\right)\mathbf{I}-\frac{\lambda}{4}\bm{\epsilon}\cdot\hat{\mathbf{z}}+\frac{\lambda^{2}}{20}\hat{\mathbf{z}}\hat{\mathbf{z}}\right]\cdot\mathbf{f}_{\alpha}.$ (15)
Here we have assumed that each sphere experiences no hydrodynamic interaction from the other, i.e., $\textbf{M}_{\alpha\beta}=\textbf{0}$, so as to focus on the higher-order effects of odd viscosity on the self-mobility of each sphere. In the above equation, one can see that the even orders of odd viscosity affect only the diagonal terms of the hydrodynamic mobility, because the diagonal contributions must be independent of the “handedness” of the fluid chirality. Moreover, the presence of anisotropy due to the odd viscosity axis $\mathbf{e}$ creates unequal forces between axial and transverse directions. Although the dimer angular dynamics can be determined by a similar way to the previous calculations, we employ another approach here by exploiting the fact that the torque-free condition is simplified in the spherical coordinates. As detained in Appendix A, the angular dynamics is found as
@@
$\dot{\phi}$ $=\frac{\lambda}{4}\frac{\dot{L}}{L}.$ (17)
It is clear from Eq. (16) that the out-of-plane dynamics of the dimmer is possible to the quadratic order or higher in $\lambda$, meaning that the odd viscosity anisotropy gives rise to the alignment behavior toward the chirality axis. Consequently, the resulting orientational dynamics shows the precessional motion as well as the reorientation of the dimer along $\mathbf{e}$, while it undergoes the monotonic expansion or contraction in its size $L(t)$. Expressions (16) and (17) are the main results of this work, where we find that the presence of chirality inside the fluid induces three-dimensional rotational oscillations depending on the expansion-contraction of the dimer. Similar alignment behavior may take place also in an anisotropic environment with a preferred direction [60]. Nevertheless, the fluid chirality we consider originates purely from the fluid chirality and the obtained alignment mechanism is distinctive, as it does not rely on external mechanism, such as ambient torque coming from the minimization of elastic energy in the case of nematic liquid crystals.
Similar to the first-order analyses, we find the orientational dynamics of a dimer vanishes, $\dot{\theta}=\dot{\phi}=0$, when $\theta=0$ and $\theta=\pi$, while for $\theta=\pi/2$, the polar-angle dynamics vanishes $(\dot{\theta}=0)$, but the evolution of $\phi$ still obeys Eq. (17). Additionally, the emergence of the quadratic $\lambda$-dependence in $\dot{\theta}$ is required from the symmetry of a force dipole as a polar vector. The combination of the polar vector and the pseudovector associated with odd viscosity then suggests the invariance under $\to-\textbf{e}$ e, from which one can deduce the even order of $\lambda$ for the polar angle evolution. The symmetry argument here is consistent with the chirotactic response of a solitary self-propelled microswimmer/active particle in chiral environments [46].
@@
while the solution to Eq. (17) is given by Eq. (14). We can see from Eq. (19) that the polar angle oscillates depending on the instantaneous ratio between $L(t)$ and $L(0)$. When the dumbbell expands and the domains move apart, i.e., $L(0)/L(t)<1$, the dimer reorients itself along the e-direction, while the contraction induces the alignment in the plane perpendicular to e without any preferred direction. This is because the systems has cylindrical symmetry about e and the two-dimensional plane $(x,y)$ is isotropic, as can be also deduced from Eq. (14).
@@
The coupling between the linear conformational change prescribed by $L(t)$ and the orientational dynamics suggests that a passive dimer undergoing thermal fluctuations may also show the rotational diffusion, which will be affected by odd viscosity. Whether it actually takes place requires an analysis of the conformational dynamics using a stochastic model. In the previous literature, there exist various different stochastic dimer models taking into account passive and active noises. Some of passive models for dimers include the Ornstein-Uhlenbeck process (OU) [61] or dimers with variable stiffness constant that is related to the recently proposed OU2 model [62]. For active dimers, on the other hand, one can consider the active version of an OU process (AOU) in the presence of active thermal fluctuations [63, 64] or dimers which are associated to molecular machines and undergo conformational oscillations through some non-equilibrium chemical process [65, 66]. To our knowledge, however, odd viscosity and the resulting asymmetric mobility tensor has not studied in the overdamped Langevin system in the context of chiral active fluids (See Ref. [67] for the discussion on underdamped systems).
In the following, we consider a passive dumbbell embedded in an odd-viscous fluid in the presence of fluctuations. The interdomain hydrodynamic interactions are again assumed to be negligible, to analyze the interplay between the anti-symmetric self-mobility and the ambient noise. To describe the dynamics in the underdamped regime, we use the following Langevin equation
$\dot{L}_{a}=\mathcal{M}_{ab}f_{b}+\xi_{a}$ (20)
with $a,b=\{\ell,\theta,\phi\}$. The first term represents the deterministic dynamics in the basis of spherical coordinates with the mobility tensor $\bm{\mathcal{M}}$, as defined in Eq. (A), and the second is the stochastic noise that satisfies $\langle\xi_{a}(t)\rangle=0$ and $\langle\xi_{a}(t)\xi_{b}(t^{\prime})\rangle=k_{\rm B}T(\mathcal{M}_{ab}+\mathcal{M}_{ba})\delta(t-t^{\prime})\equiv 2D_{ab}\delta(t-t^{\prime})$. Here $k_{\mathrm{B}}$ is the Boltzmann constant, $T$ is the temperature of the system, and $D_{ab}$ is the symmetric positive-definite diffusion tensor that depends on $\lambda$. With the choice that only the symmetric components of the diffusion tensor enter in the noise variance, it follows that $D_{ab}$ obeys a generalized Einstein relation [68, 69].
To make analytical progress we assume that $L(t)=L_{0}+\ell(t)$ where $\ell(t)\ll L_{0}$ with $L_{0}$ being the average length of the dimer and $\ell(t)$ obeying a stochastic process either continuously or discretely. This setup implies that the angular oscillation will be small in magnitude and thus similarly with the length one can assume $\theta=\theta_{0}+\delta\theta(t)$ and $\phi=\phi_{0}+\delta\phi(t)$. As a simple example, we consider the case where $\ell(t)$ follows an OU process. By utilizing Eqs. (A) and (20), the linearized Langevin dynamics are given by (see Appendix B for a derivation)
@@
where $-k\ell$ represents the force due to the conformational dynamics and the mobilities $\mu_{\ell},\mu_{\theta},$ and $\mu_{\phi}$ represent the mobilities of the corresponding velocities.
We are interested in calculating the second moment and in particular, the rotational diffusion dynamics $\langle\delta\theta^{2}\rangle+\langle\delta\phi^{2}\rangle\sim 4D^{\mathrm{eff}}_{\mathrm{r}}t$ as $t\to\infty$. By solving Eqs. (21)-(23) and performing the average over the time, the effective rotational diffusion is found as (see Appendix B for a derivation)
$D^{\mathrm{eff}}_{\mathrm{r}}$ $=\frac{1}{2L_{0}^{2}}\left(D_{\theta\theta}+\frac{D_{\phi\phi}}{\sin^{2}\theta_{0}}\right)+\frac{\mu_{\phi}^{2}}{2\mu_{\ell}L_{0}^{2}\sin^{2}\theta_{0}}$
$=D_{\mathrm{r}}+\frac{\lambda^{2}}{32L_{0}^{2}\sin^{2}\theta_{0}}\frac{a_{1}+a_{2}}{6\pi\eta^{\mathrm{e}}a_{1}a_{2}}k_{\mathrm{B}}T+\mathcal{O}(\lambda^{4}),$ (24)
where the first term $D_{\mathrm{r}}$ results from the rotational diffusion of the dimer and the second represents the odd viscosity-induced correction to it. The presence of the oddity in the system couples the orientational dynamics with the spatial conformations, leading also to the change in the rotational diffusion coefficient. In other words, the reciprocal rotational oscillations (resulting from the conformational oscillations of a dumbbell dimer) affect the diffusivity of the system. Interestingly, this chirality-induced diffusion is similar to the case where reciprocal motion influences spatial diffusion [70]. In Appendix C, we demonstrate that the observed three-dimensional dynamics may occur in odd systems in general, by analyzing the system with Lorentz forces.
@@
We have investigated the three-dimensional dynamics of a dumbbell dimer in a chiral active fluid with odd viscosity. In stark contrast to the classical case, the presence of chirality can create orthogonal forces (as a result of off-diagonal couplings in the mobility tensor), which induces motion perpendicular to the dimer axis. The oscillatory motion of a dimer gives rise to the oscillatory forces, which in turn leads to the periodic angular dynamics. To the first order in $\lambda$, we showed that only the azimuthal orientation of the dimer is affected by odd viscosity, leading to the precession dynamics around the odd viscosity axis $\mathbf{e}$. On the other hand, the dynamics of the polar angle becomes apparent in quadratic order. The obtained three-dimensional rotational dynamics provides an insight into the new odd phenomenology beyond the understanding of isotropy-compatible odd viscosity in two-dimensional experimental setups [71]. Furthermore, the observed coupling between the chirality and anisotropy would facilitate a number of new phenomenology in the analogous problems of odd diffusivity/mobility [72] in two-dimensional systems, which include the tracer dynamics [72], in many-body systems [73, 74], and polymer models in magnetic fields [75].
@@
We point out that the effective rotational dynamics obtained in Eq. (Orientational dynamics of passive dimer with thermal fluctuations) can lead to enhancement of the diffusion, depending on which model to be employed for the noise, and thus the resulting effect is model-specific. For instance, in the recently reported OU2 process, the length dynamics is affected by the stochastic dynamics of the stiffness constant [62], whereas in the AOU process, the diffusivity changes due to the presence of active thermal fluctuations [63]. Similarly in the models for chemically induced conformational oscillations, the second moment of the displacements becomes a function of the chemical reaction coordinate [66].
@@ Acknowledgements.
We thank E. Kalz for fruitful discussions. We acknowledge support from the Max Planck Center Twente for Complex Fluid Dynamics, the Max Planck Society, the Max Planck School Matter to Life, and the MaxSynBio Consortium, which are jointly funded by the Federal Ministry of Education and Research (BMBF) of Germany. Y.H. acknowledges support from JSPS Overseas Research Fellowships (Grant No. 202460086).
@@
×$\times\begin{pmatrix}1-\frac{\lambda^{2}}{10}+\frac{\lambda^{2}}{20}\cos^{2}\theta&-\frac{\lambda^{2}}{40}\sin 2\theta&-\frac{\lambda}{4}\sin\theta\\ -\frac{\lambda^{2}}{40}\sin 2\theta&1-\frac{\lambda^{2}}{10}+\frac{\lambda^{2}}{20}\sin^{2}\theta&-\frac{\lambda}{4}\cos\theta\\ \frac{\lambda}{4}\sin\theta&\frac{\lambda}{4}\cos\theta&1-\frac{\lambda^{2}}{10}\end{pmatrix}.$ (25)
The mobility tensor $\bm{\mathcal{M}}$ can be decomposed into the symmetric and antisymmetric parts, where the off-diagonal components are responsible for the angular dynamics of the dumbbell dimer. Using the equation of $\dot{L}$ to solve for $f_{n}$, and then substituting back in the equations of $\dot{\theta}$ and $\dot{\phi}$, we obtain the differential equations (16) and (17) accurate to $\mathcal{O}(\lambda^{2})$. Notice that the anisotropy imposed by the fluid chirality is reflected in the structure of the mobility tensor in Eq. (A).
Appendix B Derivation of Eqs. (21)-(23) and Eq. (Orientational dynamics of passive dimer with thermal fluctuations)
Applying the small oscillations in the spherical coordinates, such that $L(t)=L_{0}+\ell(t)$, $\theta(t)=\theta_{0}+\delta\theta(t),$ and $\phi(t)=\phi_{0}+\delta\phi(t)$, as done in the main text, we find from the Langevin equation (20) the linearized velocity vector (˙$(\dot{\ell},L_{0}\delta\dot{\theta},L_{0}\sin\theta_{0}\delta\dot{\phi})$. For an OU process, the force just obeys Hooke’s law such that $f_{n}=-k\ell$. Defining $\mu_{\ell}\equiv\mathcal{M}_{\ell\ell}$, $\mu_{\theta}\equiv\mathcal{M}_{\theta\ell},$ and $\mu_{\phi}\equiv\mathcal{M}_{\phi\ell}$ gives Eqs. (21)-(23) in the main text.
@@
Inserting this solution to the other two equations (22) and (23) for $\delta\theta$ and $\delta\phi$ gives their corresponding solutions. With these, we calculate the correlators. It is worth noting that to the second order in $\lambda$, the only correction comes from the dynamics of $\delta\phi$. More specifically, the second moments become
$\langle\delta\theta^{2}\rangle$ $\sim 2\frac{D_{\theta\theta}}{L_{0}^{2}}t,$ (27)
$\langle\delta\phi^{2}\rangle$ $\sim 2\frac{D_{\phi\phi}}{(L_{0}\sin\theta_{0})^{2}}t$
[…]
which results in Eq. (Orientational dynamics of passive dimer with thermal fluctuations). We note that $D_{\theta\theta}$ and $D_{\phi\phi}$ are smaller in the presence of odd viscosity, which resembles other odd systems, such as the suppressed rotational dynamics of a two-dimensional charged dimer in a magnetic field [75].
Appendix C Connection with the Lorentz force

### v2

@@ Introduction
We have shown that the odd viscosity, to leading order, gives rise to the precession dynamics of a dumbbell dimer about the odd viscosity direction $\hat{\mathbf{z}}$. Here we extend the model to take into account higher order contributions and investigate whether the dynamics also exhibits the out-of-plane dynamics that includes the evolution of $\theta$.
By expanding $\dot{\textbf{L}}$ in Eq. (Orientational dynamics of active dimer) up to the quadratic order in $\lambda$, the distance between the spheres evolves as
$\dot{\textbf{L}}=\frac{1}{6\pi\eta^{\rm e}}\left(\frac{1}{a_{1}}+\frac{1}{a_{2}}\right)\left[\left(1-\frac{\lambda^{2}}{10}\right)\mathbf{I}+\frac{\lambda^{2}}{20}\hat{\mathbf{z}}\hat{\mathbf{z}}-\frac{\lambda}{4}\bm{\epsilon}\cdot\hat{\mathbf{z}}\right]\cdot\mathbf{F}.$ (18)
Here we have assumed that each sphere experiences no hydrodynamic interaction from the other, i.e., $\textbf{M}_{\alpha\beta}=\textbf{0}$, so as to focus on the higher-order effects of odd viscosity on the self-mobility of each sphere. In the above equation, one can see that the even orders of odd viscosity affect only the diagonal terms of the hydrodynamic mobility, because the diagonal contributions must be independent of the “handedness” of the fluid chirality. Moreover, the presence of anisotropy due to the odd viscosity axis $\hat{\mathbf{z}}$ creates unequal forces between axial and transverse directions.
To proceed with the calculation, we employ an approach similar to the previous section, working in the frame of the spherical angular variables. As detailed in Appendix A, the angular dynamics is found as
@@
$\dot{\phi}$ $=\frac{\lambda}{4}\frac{\dot{L}}{L},$ (20)
where the latter is valid when $0<\theta<\pi$, as before. It is clear from Eq. (19) that the out-of-plane dynamics of the dimer is possible to the quadratic order or higher in $\lambda$, meaning that the odd viscosity anisotropy gives rise to the transient alignment behavior toward the chirality axis, followed by the net oscillatory motion during one cycle of the deformation. Consequently, the resulting orientational dynamics shows the precessional motion as well as the oscillatory behavior with respect to the $\hat{\mathbf{z}}-$axis, while it undergoes the monotonic expansion or contraction in its size $L(t)$. Expressions (19) and (20) are the main results of this work, where we find that the presence of chirality inside the fluid induces three-dimensional rotational oscillations depending on the expansion-contraction of the dimer. Similar alignment behavior may take place also in an anisotropic environment with a preferred direction [62]. Nevertheless, the fluid chirality we consider originates purely from the fluid chirality and the obtained alignment mechanism is distinctive, as it does not rely on external mechanism, such as ambient torque coming from the minimization of elastic energy in the case of nematic liquid crystals.
Similar to the first-order analyses, we find the orientational dynamics of a dimer vanishes, $\dot{\theta}=\dot{\phi}=0$, when $\theta=0$ and $\theta=\pi$, while for $\theta=\pi/2$, the polar-angle dynamics vanishes $(\dot{\theta}=0)$, but the evolution of $\phi$ still obeys Eq. (20). Additionally, the emergence of the quadratic $\lambda$-dependence in $\dot{\theta}$ is required from the symmetry of a force dipole as a polar vector. The combination of the polar vector and the pseudovector associated with odd viscosity then suggests the invariance under $\hat{\textbf{z}}\to-\hat{\textbf{z}}$, from which one can deduce the even order of $\lambda$ for the polar angle evolution. The symmetry argument here is consistent with the chirotactic response of a solitary self-propelled microswimmer/active particle in chiral environments [47].
@@
while the solution to Eq. (20) is given by Eq. (17). We can see from Eq. (22) that the polar angle oscillates depending on the instantaneous ratio between $L(t)$ and $L(0)$. When the dumbbell expands and the domains move apart, i.e., $L(0)/L(t)<1$, the dimer reorients itself along the $\hat{\textbf{z}}-$direction, meaning that the angle between n and the axis of odd viscosity $\hat{\textbf{z}}$ becomes smaller, while the contraction induces the the opposite effect, namely, alignment in the plane perpendicular to $\hat{\textbf{z}}$ without any preferred direction [Fig. 1(c)]. This is because the systems has cylindrical symmetry about $\hat{\textbf{z}}$ and the two-dimensional plane $(x,y)$ is isotropic, as can be also deduced from Eq. (17).
Figure 2 shows the angular dynamics of a dimer for different values of $\lambda$. The kinematic description of $L(t)$ is simply assumed to be $L(t)=L_{0}+\ell_{0}\sin(t)$, with $L_{0}$ being the average length of the dimer and $\ell_{0}$ being the maximum amplitude of oscillation. As seen in Fig. 2(a), the polar angle shows an oscillatory behavior in time, where the expansion of the dimer reorienting it towards the $z$-axis, while the contraction makes it away from the direction to align in the $(x,y)$-plane. The magnitude of the oscillation of the angle is enhanced with increasing $\lambda$, and this can be seen also for the time-evolution of $\phi$ [Fig. 2(b)].
In the above analysis, we did not take into account the hydrodynamic interaction between the spherical domains. Here we briefly examine the higher-order odd viscosity effect in the opposite regime, where the dynamics are governed solely by the hydrodynamic interaction (6), without contributions from the mobility of a sphere (5). This regime can be justified when the spheres are sufficiently small to be treated as point particles. Interestingly, from the full hydrodynamic interaction (6), we observe that the dimer does not undergo any out-of-plane dynamics due to the absence of anisotropic terms involving $\hat{\mathbf{z}}\hat{\mathbf{z}}$. This vanishing contribution to the $\theta$ dynamics holds for any value of $\lambda$, highlighting the crucial role of the particle self-mobility in the transient aligning behavior identified in Eq. (19).
In Appendix B, we discuss another type of odd systems, the system with Lorentz forces, as a parallelism to the odd-viscous system we have studied. We analyze the dynamics of a charged dimer in the presence of an external magnetic field. We demonstrate that the observed three-dimensional dynamics occur in general in other odd systems by revealing the underlying connection between chiral fluids and magnetic systems.
Figure 2: The angular dynamics of the dimer orientation as a function of time $t$ for various values of $\lambda$ when the self-mobility is taken into account up to the quadratic order in $\lambda$ and there is no hydrodynamic interaction between the domains. The dimer is initially oriented with $\theta(0)=\phi(0)=\pi/4$ and its length is prescribed as $L(t)=L_{0}+\ell_{0}\sin(t)$ for $L_{0}=1$ and $\ell_{0}=0.5$. The evolution of (a) the polar angle $\theta$ [Eq. (22)] and (b) the azimuthal angle $\phi$ [Eq. (17)].
@@
The coupling between the linear conformational change prescribed by $L(t)$ and the orientational dynamics suggests that a passive dimer undergoing thermal fluctuations may also show the rotational diffusion, which will be affected by odd viscosity. Whether it actually takes place requires an analysis of the conformational dynamics using a stochastic model. In the previous literature, there exist various different stochastic dimer models taking into account passive and active noises. Some of passive models for dimers include the Ornstein-Uhlenbeck (OU) process [63] or dimers with variable stiffness constant that is related to the recently proposed OU2 model [64]. For active dimers, on the other hand, one can consider the active version of an OU (AOU) process in the presence of active thermal fluctuations [65, 66] or dimers which are associated to molecular machines and undergo conformational oscillations through some non-equilibrium chemical process [67, 68]. To our knowledge, however, odd viscosity and the resulting asymmetric mobility tensor have not been studied in the overdamped Langevin system in the context of chiral active fluids (See Ref. [69] for the discussion on underdamped systems).
In the following, we consider a passive dumbbell embedded in an odd-viscous fluid in the presence of fluctuations. The interdomain hydrodynamic interactions are again assumed to be negligible, to analyze the interplay between the anti-symmetric self-mobility and the ambient noise. To describe the dynamics in the overdamped regime, we use the following Langevin equation
$\dot{L}_{a}=\mathcal{M}_{ab}F_{b}+\xi_{a},$ (23)
[…]
To make analytical progress we assume that $L(t)=L_{0}+\ell(t)$ where $\ell(t)\ll L_{0}$ with $L_{0}$ being the average length of the dimer and $\ell(t)$ obeying a stochastic process either continuously or discretely. This setup implies that the angular oscillation will be small in magnitude and thus similarly with the length one can assume $\theta=\theta_{0}+\delta\theta(t)$ and $\phi=\phi_{0}+\delta\phi(t)$. As a simple example, we consider the case where $\ell(t)$ follows an OU process. By utilizing Eqs. (23) and (A), the linearized Langevin dynamics are given by (see Appendix C for a derivation)
@@
where $k\ell$ represents the force due to the conformational dynamics and the mobilities $\mu_{\ell},\mu_{\theta},$ and $\mu_{\phi}$ represent the mobilities of the corresponding velocities. It is worth noting that the angular diffusivities, $D_{\theta\theta}$, and $D_{\phi\phi}$, are not related to $\mu_{\theta}$ and $\mu_{\phi}$. In particular, as shown in Appendices A-C, $\mu_{\theta}=\mathcal{M}_{\theta\ell}$ and $\mu_{\phi}=\mathcal{M}_{\phi\ell}$, where for the noise strength we have, $D_{\theta\theta}=k_{\rm B}T\mathcal{M}_{\theta\theta}$ and $D_{\phi\phi}=k_{\rm B}T\mathcal{M}_{\phi\phi}$.
We are interested in calculating the second moment, particularly the angular mean squared displacement given by $\langle\delta\theta^{2}(t)\rangle+\langle\delta\phi^{2}(t)\rangle$ at the early dynamics of $\mu_{\ell}kt\ll 1$, where our linearization scheme is valid. By solving Eqs. (24)-(26) and calculating the corresponding correlators, the angular mean squared displacement is found as (see Appendix C for a derivation)
$\langle\delta\theta^{2}(t)\rangle+\langle\delta\phi^{2}(t)\rangle$
$\simeq\frac{2}{L_{0}^{2}}\left(D_{\theta\theta}+\frac{D_{\phi\phi}}{\sin^{2}\theta_{0}}\right)t+k_{\mathrm{B}}T\frac{a_{1}+a_{2}}{6\pi\eta^{\mathrm{e}}a_{1}a_{2}}\frac{\lambda^{2}k^{2}}{24L_{0}^{2}}t^{3},$ (27)
where the first term results from the rotational diffusion of the dimer and the second represents the odd viscosity-induced correction to it, which is cubic in time. The presence of the oddity in the system couples the orientational dynamics with the spatial conformations, leading also to the change in the rotational diffusion coefficient. In other words, the rotational oscillations (resulting from the reciprocal conformational changes of a dumbbell dimer) affect the diffusivity of the system. Interestingly, this chirality-induced diffusion is similar to the case where reciprocal motion influences spatial diffusion [72].
@@
We have investigated the three-dimensional dynamics of a dumbbell dimer in a chiral active fluid with odd viscosity. In stark contrast to the classical case, the presence of chirality can create orthogonal forces (as a result of off-diagonal couplings in the mobility tensor), which induces motion perpendicular to the dimer axis. The oscillatory motion of a dimer gives rise to the oscillatory forces, which in turn leads to the periodic angular dynamics. To the first order in $\lambda$, we showed that only the azimuthal orientation of the dimer is affected by odd viscosity, leading to the precession dynamics around the odd viscosity axis $\hat{\textbf{z}}$. On the other hand, the dynamics of the polar angle becomes apparent in quadratic order. The obtained three-dimensional rotational dynamics provides an insight into the new odd phenomenology beyond the understanding of isotropy-compatible odd viscosity in two-dimensional experimental setups [73]. Furthermore, the observed coupling between the chirality and anisotropy would facilitate a number of new phenomenology in the analogous problems of odd diffusivity/mobility [74] in two-dimensional systems, which include the tracer dynamics [74], in many-body systems [75, 76], and polymer models in magnetic fields [77].
In principle, modifications to the dimer dynamics are expected when higher-order contributions beyond the leading hydrodynamic interaction [Eq. (6)] are considered. One possible extension is the Rotne–Prager–Yamakawa tensor, which includes the next higher-order term in the far-field approximation [78, 79]. However, extensions that incorporate both higher-order corrections and arbitrary odd viscosity values are rarely reported. Although deriving a Rotne–Prager–Yamakawa-type hydrodynamic mobility for odd fluids is beyond the scope of this work, such corrections could provide valuable insights into dimer dynamics, particularly regarding the influence of dimer domains on the hydrodynamic interactions.
@@
We point out that studying full non-linear rotational dynamics can possibly lead to enhancement of the diffusion, depending on which model to be employed for describing the spatial oscillations, and thus the resulting effect is model-specific. For instance, in the recently reported OU2 process, the length dynamics is affected by the stochastic dynamics of the stiffness constant [64], whereas in the AOU process, the diffusivity changes due to the presence of active thermal fluctuations [65]. Similarly in the models for chemically induced conformational oscillations, the second moment of the displacements becomes a function of the chemical reaction coordinate [68].
@@ Acknowledgements.
We thank E. Kalz and A. Vilfan for fruitful discussions. We acknowledge support from the Max Planck Center Twente for Complex Fluid Dynamics, the Max Planck Society, the Max Planck School Matter to Life, and the MaxSynBio Consortium, which are jointly funded by the Federal Ministry of Education and Research (BMBF) of Germany. Y.H. acknowledges support from JSPS Overseas Research Fellowships (Grant No. 202460086).
@@
$\times\begin{pmatrix}1-\frac{\lambda^{2}}{10}+\frac{\lambda^{2}}{20}\cos^{2}\theta&-\frac{\lambda^{2}}{40}\sin 2\theta&-\frac{\lambda}{4}\sin\theta\\ -\frac{\lambda^{2}}{40}\sin 2\theta&1-\frac{\lambda^{2}}{10}+\frac{\lambda^{2}}{20}\sin^{2}\theta&-\frac{\lambda}{4}\cos\theta\\ \frac{\lambda}{4}\sin\theta&\frac{\lambda}{4}\cos\theta&1-\frac{\lambda^{2}}{10}\end{pmatrix}$. (28)
@@
Appendix B Connection with the Lorentz force